Simplify the expression by using a Double-Angle Formula or a Half-Angle Formula.
(2 tan(5θ))/ (1 − tan^2(5θ))

Answer by ikleyn(52794)

.

(2 tan(5θ))/ (1 − tan^2(5θ)) = tan(5θ + 5θ) = tan(10θ).    ANSWER